to cover a certain distance a passenger jeep traveling at 48km per hour requires 3 hours longer than a station wagon traveling at 72km per hour. find the distance.
Answer by ccs2011(207)

Let t be the time it takes the station wagon to travel distance d.
d = r*t
d= 72t
d= 48(t+3)
using substitution
72t = 48(t+3)
72t = 48t + 144
24t = 144
t = 6
Substitute this value in to find d
d = 72*6
d = 432
Therefore the distance traveled is 432 km